# A Unified Representation of Density-Power-Based Divergences Reducible to M-Estimation

### v1

@@ II-C Density-power-based divergences with functional flexibility
PS-type Bridge Density Power (15) $\frac{1}{\lambda_{2}}\left[\left(\lambda_{1}+\lambda_{2}z^{1+\gamma}\right)^{\frac{1}{1+\gamma}}-\lambda_{1}^{\frac{1}{1+\gamma}}\right]$ $\lambda_{1}\geq 0,\lambda_{2}\geq 0$ $\lambda_{1}=0$: scaled PS
(PSBDP) [8, 9] $\lambda_{2}\to 0$: scaled DP
Combination of BH and PSBDP (17) $\frac{1}{\lambda_{2}}\left[\left(\lambda_{1}+\lambda_{2}z^{1+\gamma}\right)^{\frac{\kappa}{1+\gamma}}-\lambda_{1}^{\frac{\kappa}{1+\gamma}}\right]$ $\lambda_{1}\geq 0,\lambda_{2}\geq 0,\kappa\geq 1$ $\lambda_{1}=0,\lambda_{2}\neq 0$: scaled BH
@@ III-A Existing divergences
Previous studies [8, 9] have not shown that the BDPD is a special case of the Bregman divergence in the sense of Definition 2. Previous studies [8] [^1] In [8], $\bm{\lambda}$ is constrained by $\lambda_{1}+\lambda_{2}=1$., [9] have also pointed out that log-type BDPD can be reduced to log-$\gamma$-divergence [4] and DPD [1], but only when ($\lambda_{1}=0,\lambda_{2}=1$) and ($\lambda_{1}=1,\lambda_{2}\to 0$). In practice, even in the case of the log-type BDPD, it reduces to the log-$\gamma$-divergence [4] when ($\lambda_{1}=0,\lambda_{2}\neq 0$) and to the DPD [1] when ($\lambda_{1}\neq 0,\lambda_{2}\to 0$) (see Appendix A).
@@ Appendix B Hölder divergence [10]
When we set $H(z)=-\kappa^{\frac{1+\gamma}{\kappa}}|z-1+1/\kappa|^{\frac{1+\gamma}{\kappa}}{\rm sign}(z-1+1/\kappa),\;k\geq 1$, the HCE (22) reduces to the transformation of the BHCE (14) as follows:
@@ Appendix C Special cases of Norm-based Bregman density power divergence
We provide the specific cross-entropy of NB-DPCE (10) discussed in Section III-B. We also restate the cross-entropy (17) that connects BHCE (14) and BDCE (15) as follows:
@@
$D_{\bm{\lambda},\kappa,\gamma}(q,p)=\frac{\kappa}{\lambda_{2}\gamma}\Bigl{(}\lambda_{1}+\lambda_{2}\bigl{\langle}p^{1+\gamma}\bigr{\rangle}\Bigr{)}^{\frac{\kappa}{1+\gamma}}\left[1-\frac{1}{\kappa}-\frac{\lambda_{1}+\lambda_{2}\bigl{\langle}qp^{\gamma}\bigr{\rangle}}{\lambda_{1}+\lambda_{2}\bigl{\langle}p^{1+\gamma}\bigr{\rangle}}\right]+\frac{\kappa}{\lambda_{2}\gamma}\Bigl{(}\lambda_{1}+\lambda_{2}\bigl{\langle}p^{1+\gamma}\bigr{\rangle}\Bigr{)}^{\frac{\kappa}{1+\gamma}}.$
@@
When $\lambda_{1}\neq 0,\lambda_{2}\neq 0,\kappa$, (23) reduces to the PS-type BDCE (15) as follows:
@@ Appendix E Proof of Theorem 3
$\lim_{x_{o}\to\infty}\psi(x_{o},\theta)=-\left\langle p_{\theta}^{1+\gamma}\right\rangle\left\langle p_{\theta}^{1+\gamma}s_{\theta}\right\rangle\left\|p_{\theta}\right\|_{1+\gamma}\phi^{\prime\prime}_{\gamma}\left(\left\|p_{\theta}\right\|_{1+\gamma}\right)=0$

### v2

@@ II-C Density-power-based divergences with functional flexibility
PS-type Bridge Density Power (15) $\frac{1}{\lambda_{2}}\left[\left(\lambda_{1}+\lambda_{2}z^{1+\gamma}\right)^{\frac{1}{1+\gamma}}-\lambda_{1}^{\frac{1}{1+\gamma}}\right]$ $\lambda_{1}\geq 0,\lambda_{2}\geq 0$ $\lambda_{1}=0,\lambda_{2}\neq 0$: scaled PS
(PSBDP) [8, 9] $\lambda_{1}\neq 0,\lambda_{2}\to 0$: scaled DP
Combination of BH and PSBDP (17) $\frac{1}{\lambda_{2}}\left[\left(\lambda_{1}+\lambda_{2}z^{1+\gamma}\right)^{\frac{\kappa}{1+\gamma}}-\lambda_{1}^{\frac{\kappa}{1+\gamma}}\right]$ $\lambda_{1}\geq 0,\lambda_{2}\geq 0,$ $\lambda_{1}=0,\lambda_{2}\neq 0$: scaled BH
@@ III-A Existing divergences
Previous studies [8, 9] have not shown that the BDPD is a special case of the Bregman divergence in the sense of Definition 2. Previous studies [8] [^1] In [8], $\bm{\lambda}$ is constrained by $\lambda_{1}+\lambda_{2}=1$., [9] have also pointed out that log-type BDPD can be reduced to log-$\gamma$-divergence [4] and DPD [1], but only when ($\lambda_{1}=0,\lambda_{2}=1$) and ($\lambda_{1}=1,\lambda_{2}\to 0$). In practice, even in the case of the log-type BDPD, it reduces to the log-$\gamma$-divergence [4] when ($\lambda_{1}=0,\lambda_{2}\neq 0$) and to the DPD [1] when ($\lambda_{1}\neq 0,\lambda_{2}\to 0$) (see Appendix A, Example 8).
@@ Appendix B Hölder divergence [10]
When we set $H(z)=-\kappa^{\frac{1+\gamma}{\kappa}}|z-1+1/\kappa|^{\frac{1+\gamma}{\kappa}}{\rm sign}(z-1+1/\kappa),\;\kappa\geq 1$, the HCE (22) reduces to the transformation of the BHCE (14) as follows:
@@ Appendix C Special cases of Norm-based Bregman density power divergence
We provide the specific cross-entropy of NB-DPCE (10) discussed in Section III-B. We also restate the cross-entropy (17) that connects BHCE (14) and BDPCE (15) as follows:
@@
$D_{\bm{\lambda},\kappa,\gamma}(q,p)=\frac{\kappa}{\lambda_{2}\gamma}\Bigl{(}\lambda_{1}+\lambda_{2}\bigl{\langle}p^{1+\gamma}\bigr{\rangle}\Bigr{)}^{\frac{\kappa}{1+\gamma}}\left[1-\frac{1}{\kappa}-\frac{\lambda_{1}+\lambda_{2}\bigl{\langle}qp^{\gamma}\bigr{\rangle}}{\lambda_{1}+\lambda_{2}\bigl{\langle}p^{1+\gamma}\bigr{\rangle}}\right]+\frac{1}{\lambda_{2}\gamma}\Bigl{(}\lambda_{1}+\lambda_{2}\bigl{\langle}q^{1+\gamma}\bigr{\rangle}\Bigr{)}^{\frac{\kappa}{1+\gamma}}.$
@@
When $\lambda_{1}\neq 0,\lambda_{2}\neq 0,\kappa=1$, (23) reduces to the PS-type BDPCE (15) as follows:
@@
When, $\lambda_{1}=0,\lambda_{2}\neq 0,\kappa=1$, (23) reduces to a constant multiple of the PSCE (6) as follows:
$d_{\bm{\lambda},\kappa,\gamma}(q,p)=\lambda_{2}^{-\frac{\gamma}{1+\gamma}}d_{\gamma}^{\rm PS}(q,p).$
@@ Appendix E Proof of Theorem 3
$\lim_{x_{o}\to\infty}\psi(x_{o},\theta)=\left\langle p_{\theta}^{1+\gamma}\right\rangle\left\langle p_{\theta}^{1+\gamma}s_{\theta}\right\rangle\left\|p_{\theta}\right\|_{1+\gamma}\phi^{\prime\prime}_{\gamma}\left(\left\|p_{\theta}\right\|_{1+\gamma}\right)=0$